24. SELECTION oF PUBLic HOUSING APPLICANTS The City Housing Authority has received 50 applications from qualified applicants for eight low-income apartments. Three of the apartments are on the north side of town, and five are on the south side. If the apartments are to be assigned by means of a lottery, what is the probability that: a. A specific qualified applicant will be selected for one of these apartments? b. Two specific qualified applicants will be selected for apartments on the same side of town?
22. QUALITY CONTROL Electronic baseball games manufactured by Tempco Electronics are shipped in lots of 24. Before shipping, a quality-control inspector randomly selects a sample of 8 from each lot for testing. If the sample con- tains any defective games, the entire lot is rejected. What is the probability that a lot containing exactly 2 defective games will still be shipped?
20. A customer at Cavallaro's Fruit Stand picks a sample of 3 oranges at random from a crate containing 60 oranges, of which 4 are rotten. What is the probability that the sample contains 1 or more rotten oranges?
18. PERSONNEL SELECTION Jacobs & Johnson, an accounting firm, employs 14 accountants, of whom 8 are CPAs. If a delegation of 3 accountants is randomly selected from the firm to attend a conference, what is the probability that 3 CPAs will be selected?
Answer #1

24)

a. For a particular qualified applicant to be selected, number of ways = 49C7 (As given he is chosen, the other 7 winners will be chosen from 49 remaining applicants)

Total number of ways of selecting winners = 50C8

Probability = 49C7 / 50C8 = 8/50 =0.16

b)
Two applicants will be selected from the same side of town
= P(Two will be selected from North) + P(Two will be selected from south)
= = 3/50 x 2/49 + 5/50x4/49
=0.01061